OpenStudy (anonymous):

Please answet :) Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. y = x2 + 3x + 2 y = x2 + 2x + 1 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1?

OpenStudy (anonymous):

help please!

OpenStudy (campbell_st):

the 1st quadratic factorise to see where the curve cuts the x-axis (x +2)(x +1) = 0 will result in x = -2 and x - 1 this means the parabola \[y = x^2 + 3x + 2\] is concave up, cuts the x-axis at x = -2 and x = -1 has a line of symmetry at x = - 1.5 has a minimum value of y = -0.25 cuts the y axis at y = 2